OpenStudy (anonymous):

WILL GIVE METAL!!!! A cylinder has a radius of 5x+2 and a height of 2x+8. Which polynomial in standard form best describes the total volume of the cylinder? Use the formula V=TT r^2h for the volume of a cylinder. *100TTx^4+880TTx^3+2256TTx+256TT *50TTx^3+240TTx^2+168TTx+32TT *20TTx^3+168TTx^2+384TTx+128TT *10TT^2+44TTx+16TT TT=pie sign

OpenStudy (anonymous):

@zepdrix HELP PLEASE.

OpenStudy (anonymous):

@Nnesha PLEASE HELP.

OpenStudy (anonymous):

@kropot72 HELP?

Nnesha (nnesha):

first replace r by 5x+2

OpenStudy (anonymous):

=4

Nnesha (nnesha):

\[\huge\rm V =(5x +2)^2 (2x +8) \pi\] do

Nnesha (nnesha):

wait i have to login to another lattop

OpenStudy (anonymous):

0?

Nnesha (nnesha):

0 what ??

OpenStudy (anonymous):

I put in the equation into my calculator and it said 0?

Nnesha (nnesha):

which calculator ??

OpenStudy (anonymous):

a graphing calculator

Nnesha (nnesha):

okay but you have to foil it (5x + 2) ^2 which is same as (5x+2) times (5x+2)

OpenStudy (anonymous):

100.53

Nnesha (nnesha):

if you look at ur answer choices there are VARIABLES "X" this is not "multiplication" sign it is x so how did you get 100.53 ??? answer suppose to be with variable x

OpenStudy (anonymous):

I typed (5x+2)(5x+2)(2x+8)TT into the calculator

Nnesha (nnesha):

ti84 calculator ?

OpenStudy (anonymous):

idk one that my school told me to download Wabbitemu is what it is called

Nnesha (nnesha):

okay give me a sec to look at that

Nnesha (nnesha):

does it have yellow button ??

OpenStudy (anonymous):

its Texas Instruments TI-83 Plus Silver Edition

Nnesha (nnesha):

ohh yeah so that's not gonna work for variables :(

Nnesha (nnesha):

do you know foil method ??

OpenStudy (anonymous):

kind of.

Nnesha (nnesha):

okay lets do it together (5x+2)(5x+2) do it first and let me know what you get

Nnesha (nnesha):

foil (5x+2)(5x+2)

OpenStudy (anonymous):

25x+20x+4?

Nnesha (nnesha):

yep right but 5x times 5x = ??

Nnesha (nnesha):

when you multiply variables you should add their exponent :)

OpenStudy (anonymous):

25x^2

Nnesha (nnesha):

yep right now \[\huge\rm (25x^2 + 20x + 4)(2x+8)\] again foil method

OpenStudy (anonymous):

do I multiply 25x^2 times 20x+4 or 2x+8?

Nnesha (nnesha):

nope you have to multiply 2x and 8 with everything in this parentheses ( 25x^2 +20x+4)

Nnesha (nnesha):

\[\huge\rm 2x(25x^2 + 20x + 4) \] \[\huge\rm 8(25x^2 + 20x + 4) \]

OpenStudy (anonymous):

I know this is probably wrong be I tried 50x^2+40x+8x 200x^2+160x+32

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

again 2x times 25x^2 = ??

Nnesha (nnesha):

\[\huge\rm x^n \times x^m = x^{n+m}\]

OpenStudy (anonymous):

I honestly don't know..i have been working on this problem for over an hour and I'm completely done. thank you for all you time and help but I will ask for help on Monday.

Nnesha (nnesha):

well almost done i'm trying to say that when you multiply variables you should add their exponent so 2x^1 times 25x^2 = 50x^{1+2} = ?

OpenStudy (anonymous):

ok then what?

Nnesha (nnesha):

after that just multiply everything by pi

OpenStudy (anonymous):

3.14? @Nnesha

Nnesha (nnesha):

nope just pi not pi value

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @L<3VE I know this is probably wrong be I tried 50x^2+40x+8x 200x^2+160x+32 \(\color{blue}{\text{End of Quote}}\) please correct his mistake 2x times 25x^2 = ?? 2x times 20x = ??

Nnesha (nnesha):

this**

OpenStudy (anonymous):

so the answer is a?

Nnesha (nnesha):

well 2x times 25x^2 = ?? this is suppose to be first term

OpenStudy (anonymous):

50^2 which is B

Nnesha (nnesha):

x^1 times x^2 = ??? :(

OpenStudy (anonymous):

Sorry 50TTx^3

Nnesha (nnesha):

\(\huge\color{green}{{\rm yes}!}\) and yes b

OpenStudy (anonymous):

Thank you...

Nnesha (nnesha):

np :)